# Unraveling time-varying causal effects of multiple exposures: integrating Functional Data Analysis with Multivariable Mendelian Randomization

### v2

@@ 5 Discussion
MV-FMR addresses important methodological gaps in the growing literature on longitudinal MR. Previous approaches have considered functional methods for single exposures [cao2016timevarying, Tian2024] or multivariable methods with discretized time [sanderson2022timevarying]. In contrast, our framework jointly models multiple exposure processes as continuous functions of time. It leverages FPCA to represent longitudinal exposures, accommodating sparse and irregular measurements, and implements a data-driven cross-validation procedure for selecting the optimal number of components. By explicitly accounting for overlapping genetic instruments, MV-FMR allows direct and indirect effects to be disentangled even in the presence of pleiotropy. The framework provides estimation procedures for both continuous and binary outcomes, with theoretical justification for identifiability under standard multivariable MR assumptions.
Extensive simulation studies demonstrated the robustness and accuracy of MV-FMR across a wide range of scenarios, including nonlinear effects, horizontal pleiotropy, and mediation. The method reliably recovered the true functional forms of causal effects and consistently outperformed univariable functional MR approaches. The use of a robust estimation procedure ensured that weak instruments do not bias causal effect estimates. We further conducted sensitivity analyses varying instrument overlap, mediation strength, exposure–outcome functional forms, and data sparsity, demonstrating the method’s robustness under diverse practical conditions.
@@
The practical relevance of the proposed MV-FMR framework was demonstrated through an application to UK Biobank data, examining the time-varying causal effects of SBP and BMI on CAD risk. This analysis was intended primarily as a proof-of-concept, demonstrating how the proposed framework can recover well-established causal structures while extending them to a functional setting. The results highlighted ages 50–60 as a period during which both SBP and BMI exerted their strongest estimated effects, with attenuated effects at older ages. The comparison between multivariable and univariable functional MR provides evidence consistent with established mediation mechanisms. While SBP estimates were quite unchanged between univariable and multivariable models, BMI effects were attenuated in the joint analysis, indicating that part of BMI’s association with CAD is captured through blood pressure pathways. This pattern mirrors findings from traditional cross-sectional MV-MR, where partial mediation of the BMI-CAD relationship by SBP has been reported and where univariable and multivariable estimates similarly exhibit overlapping confidence intervals [Gill2021]. Taken together, these findings indicate that MV-FMR behaves as expected in the presence of mediation, while providing a functional representation of effects across age.
@@
Several limitations should be acknowledged. MV-FMR requires longitudinal measurements for each exposure within the same age range, which may not be available in all datasets. The computational burden increases with the number of exposures, as functional modeling is more demanding than cross-sectional analysis. Additionally, measurement error in exposures or any violations of the core MR assumptions could bias the causal estimates. While our simulation studies indicate that MV-FMR performs well even with moderately weak instruments, increasing the complexity of the underlying time-varying effects can reduce instrument strength. As the number of functional components used to represent exposures grows, the effective conditional F-statistics can decrease, potentially introducing bias if they fall below a critical threshold. Accordingly, tailored simulation analyses are essential to evaluate whether the available instruments and data structure are sufficient to support reliable inference in a given application. Such pre-analysis assessment allows researchers to evaluate model feasibility and anticipate potential limitations in estimating causal effects. Finally, although our empirical application demonstrates the method’s feasibility, it was primarily intended as a proof of concept; broader applications are necessary to fully establish practical utility across diverse epidemiological contexts, and the choice and timing of outcome measurement relative to exposures should be carefully examined through extensive sensitivity analyses.

### v3

@@ 5 Discussion
MV-FMR addresses important methodological gaps in the growing literature on longitudinal MR. Previous approaches have considered functional methods for single exposures [2, 29] or multivariable methods with discretized time [20]. In contrast, our framework jointly models multiple exposure processes as continuous functions of time. It leverages FPCA to represent longitudinal exposures, accommodating sparse and irregular measurements, and implements a data-driven cross-validation procedure for selecting the optimal number of components. By explicitly accounting for overlapping genetic instruments, MV-FMR allows direct and indirect effects to be disentangled even in the presence of pleiotropy. The framework provides estimation procedures for both continuous and binary outcomes, with theoretical justification for identifiability under standard multivariable MR assumptions.
@@
Extensive simulation studies demonstrated the robustness and accuracy of MV-FMR across a wide range of scenarios, including nonlinear effects, horizontal pleiotropy, and mediation. The method reliably recovered the true functional forms of causal effects and consistently outperformed univariable functional MR approaches. The use of a robust estimation procedure ensured that weak instruments do not bias causal effect estimates. We further conducted sensitivity analyses varying instrument overlap, mediation strength, exposure-outcome functional forms, and data sparsity, demonstrating the method’s robustness under diverse practical conditions. The practical relevance of the proposed MV-FMR framework was demonstrated through an application to UK Biobank data, examining the time-varying causal effects of SBP and BMI on CAD risk. This analysis was intended primarily as a proof-of-concept, demonstrating how the proposed framework can recover well-established causal structures while extending them to a functional setting. The results highlighted ages 50-60 as a period during which both SBP and BMI exerted their strongest estimated effects, with attenuated effects at older ages. The comparison between multivariable and univariable functional MR provides evidence consistent with established mediation mechanisms. While SBP estimates were quite unchanged between univariable and multivariable models, BMI effects were attenuated in the joint analysis, indicating that part of BMI’s association with CAD is captured through blood pressure pathways. This pattern mirrors findings from traditional cross-sectional MV-MR, where partial mediation of the BMI-CAD relationship by SBP has been reported and where univariable and multivariable estimates similarly exhibit overlapping confidence intervals [5]. Taken together, these findings indicate that MV-FMR behaves as expected in the presence of mediation, while providing a functional representation of effects across age.
@@
Several limitations should be acknowledged. MV-FMR requires longitudinal measurements for each exposure within the same age range, which may not be available in all datasets. The computational burden increases with the number of exposures, as functional modeling is more demanding than cross-sectional analysis. Additionally, measurement error in exposures or any violations of the core MR assumptions could bias the causal estimates. While our simulation studies indicate that MV-FMR performs well even with moderately weak instruments, increasing the complexity of the underlying time-varying effects can reduce instrument strength. As the number of functional components used to represent exposures grows, the effective conditional F-statistics can decrease, potentially introducing bias if they fall below a critical threshold. Accordingly, tailored simulation analyses are essential to evaluate whether the available instruments and data structure are sufficient to support reliable inference in a given application. Such pre-analysis assessment allows researchers to evaluate model feasibility and anticipate potential limitations in estimating causal effects. Finally, although our empirical application demonstrates the method’s feasibility, it was primarily intended as a proof of concept. For instance, our analysis conditioned on survival to age 50 without prevalent CAD, potentially selecting for a cardiovascularly resilient subset of the population. As high-risk individuals develop CAD early and are excluded from the baseline, the remaining cohort becomes progressively enriched with individuals resistant to the atherogenic effects of adiposity. While MR mitigates traditional confounding, it cannot entirely eliminate this form of survival bias, which may attenuate the estimated causal effects. Consequently, broader applications are necessary to fully establish practical utility across diverse epidemiological contexts, and the choice and timing of outcome measurement relative to exposures should be carefully examined through extensive sensitivity analyses.